OpenStudy (anonymous):

McCleary's Office Supply Store is having a sale on desk chairs. All chairs are 33% off the original price. The original price of one chair is $183.00. What is the sale price?

OpenStudy (anonymous):

basically the question is asking what is 33% of 183 (to find your discount) and then you subtract that answer from the original price of 183 to get your sale price. " of " means multiply turn your percent to a decimal --- 33% = 0.33 now your problem is : 0.33 * 183 = 60.39 183 - 60.39 = 122.61 <== sale price Another way.. if the sale price is 33% off of 183, that means you are paying 67% of 183 67% of 183 = .67 * 183 = 122.61 you will arrive at the same answer either way you do it.